OpenStudy (anonymous):

Solve the following using the quadratic formula. Round to the nearest tenth if necessary. 2x2 + 9x + 4 = 0 A. {-4, -0.5} B. {-16, -2} C. {-8, -1} D. no real solution

OpenStudy (anonymous):

if you see that the answers are rational numbers, try factoring first. no need for formula

OpenStudy (anonymous):

try \[(x+4) (2 x+1) = 0\]

OpenStudy (anonymous):

I'm so confused, I don't get what to do.

OpenStudy (anonymous):

Help someone please ?

jhonyy9 (jhonyy9):

x1,2=(-9+/- sqrt(81-32))/4 =(-9+/- 7)/4 = -4 and -2/4 = -1/2 solutions are -4 and -1/2

OpenStudy (anonymous):

Thank you! :)

jhonyy9 (jhonyy9):

yw